A man borrows 850 taka from bank. The bill is renewed after every half year at an increase of 21% . After how much time amount will become 8500 taka.


1
Expert's answer
2021-01-21T19:34:56-0500

"FV=PV(1+i)^n"

where "FV" is future value, "PV" is present value, "i" is the interest per compounding period, "n" is the number of compounding periods.

"n=\frac{lnFV-lnPV}{ln(1+i)}"


"n=\frac{ln8500-ln850}{ln1.21}=12.08=6.04" years